# Accelerating Greedy Coordinate Gradient and General Prompt Optimization via Probe Sampling

### v2

@@ Abstract
Safety of Large Language Models (LLMs) has become a critical issue given their rapid progresses. Greedy Coordinate Gradient (GCG) is shown to be effective in constructing adversarial prompts to break the aligned LLMs, but optimization of GCG is time-consuming. To reduce the time cost of GCG and enable more comprehensive studies of LLM safety, in this work, we study a new algorithm called Probe sampling. At the core of the algorithm is a mechanism that dynamically determines how similar a smaller draft model’s predictions are to the target model’s predictions for prompt candidates. When the target model is similar to the draft model, we rely heavily on the draft model to filter out a large number of potential prompt candidates. Probe sampling achieves up to $5.6$ times speedup using Llama2-7b-chat and leads to equal or improved attack success rate (ASR) on the AdvBench. Furthermore, probe sampling is also able to accelerate other prompt optimization techniques and adversarial methods, leading to acceleration of 1.8×1.8$\times$1.8 × for AutoPrompt, 2.4×2.4$\times$2.4 × for APE and 2.4×2.4$\times$2.4 × for AutoDAN.
@@ 1 Introduction
Ensuring the safety of Large Language Models (LLMs)  (Brown et al., 2020; Chowdhery et al., 2022; Touvron et al., 2023; Jiang et al., 2023) has become a central theme of research. Despite continuous efforts, LLMs are prone to generate objectionable contents in various scenarios including using an adversarial suffix (Zou et al., 2023), further finetuning (Qi et al., 2023; Lermen et al., 2023), ciphering (Yuan et al., 2023) and multilingual settings (Deng et al., 2023). Among effective LLM adversarial attack works, Greedy Coordinate Gradient (GCG) (Zou et al., 2023) present a general and universal method as briefly illustrated in Figure 1.
@@
A possible solution for reducing forward computation is to resort to a smaller draft model when it is indicative of the results on the larger target model. This intuition has been applied in speculative sampling (Chen et al., 2023; Leviathan et al., 2023) for decoding. In speculative sampling, the target model acts as a verifier that accepts or rejects the decoded tokens. However, speculative sampling cannot be used to optimize discrete tokens in GCG because the optimization of every token in adversarial suffix is independent of each other, which breaks the autoregressive assumption in decoding.
Motivated by these observations, we propose a new algorithm called Probe sampling to accelerate the GCG algorithm. Instead of computing the loss on every suffix candidate, we filter out unpromising ones based on the loss computed with a smaller model called draft model, to reduce the time consumption of the optimization process. Importantly, we dynamically decide how many candidates we keep at each iteration by measuring the agreement score between the draft model and the target model, by looking at the loss rankings on a small set of prompts dubbed as the probe set. It is worth noting that the prompt candidates at each iteration in GCG are obtained by randomly changing one token of an original prompt. As a result, the agreement score is adaptive to the original prompt. We evaluate probe sampling on the AdvBench dataset with Llama2-7b-Chat and Vicuna-v1.3 as the target models and a significantly smaller model GPT-2 Radford et al. (2019) as the draft model. Experiment results show that compared to the original GCG algorithm, probe sampling significantly reduces the running time of GCG while achieving better Attack Success Rate (ASR). Specifically, with Llama2-7b-Chat, probe sampling achieves $3.5$ times speedup and an improved ASR of $81.0$ compared to GCG with $69.0$ ASR. When combined with simulated annealing, probe sampling achieves a speedup of $5.6$ times with a better ASR of $74.0$.
Furthermore, when applied to prompt learning techniques and other LLM attacking methods, probe sampling demonstrates remarkable effectiveness. Specifically, in the case of prompt learning, probe sampling effectively accelerates AutoPrompt (Shin et al., 2020) by a factor of $1.8$. Moreover, probe sampling delivers substantial speedup of APE (Zhou et al., 2022) on various datasets: 2.3×2.3$\times$2.3 × on GSM8K, 1.8×1.8$\times$1.8 × on MMLU and 3.0×3.0$\times$3.0 × on BBH. In the case of other attacking method such as AutoDAN (Liu et al., 2023a), probe sampling achieve a speedup of 2.3×2.3$\times$2.3 × and 2.5×2.5$\times$2.5 × on AutoDAN-GA and AutoDAN-HGA respectively.
@@ 2.1 Background: Greedy Coordinate Gradient
where $x$ is a prompt that contains a harmful user query such as “Tell me how to build a bomb", $y$ is the target sentence “Sure, here is how to build a bomb", and $s$ is the adversarial suffix that is optimized to induce the generation of $y$. $p$ is the probability of a sentence output by a LLM. This objective can be decomposed into the summation of the negative log likelihood of individual tokens in the target sentence like a typical language modeling objective. $s$ is set to be a fixed length string in the GCG algorithm.
The optimization of the adversarial suffix $s$ is a non-trivial problem. Prior works Guo et al. (2021); Wen et al. (2023) based on Gumbel-Softmax Jang et al. (2016); Maddison et al. (2016) and soft prompt tuning Lester et al. (2021) have achieved limited success, probably because the LLMs are well-aligned and the exceptionally large models magnifies the difference between a discrete choice and its continuous relaxations.
@@ 2.2 Probe Sampling
After obtaining the agreement score, we keep $(1-\alpha)*B/R$ candidates where $(1-\alpha)*B$ means that the filtered set size is a scale-down of the previous batch size $B$ and $R$ is a hyperparameter that determines a further scale down. When $\alpha$ is close to 00, meaning little agreement between the two models, we will use a filtered set size of $B/R$. When $\alpha$ goes to $1$, we almost filter out most of the candidates.
With the filtered size determined, we can readily rank the candidates according to the draft model and filter the ones with higher losses. Finally, we evaluate the final loss on the filtered set using the target model and select the best candidate.
@@ 2.3 Applying Probe Sampling to Other Prompt Optimization Methods
Although prompt sampling was designed to accelerate GCG, the general idea of reducing forward computation can be applied on other prompt optimization methods, where there is usually a process of sampling prompt candidates and evaluating their performances. To see whether probe sampling can effectively accelerate other methods, we also apply probe sampling to two prompt learning methods AutoPrompt (Shin et al., 2020) and APE (Zhou et al., 2022). In addition, we apply probe sampling on AutoDAN (Liu et al., 2023a), a genetic algorithm that can find natural jailbreak prompts.
@@ 3.1 Experiment Details
Following Zou et al. (2023), we use Attack Success Rate (ASR) as the evaluation metric for GCG and AutoDAN, which is defined as the percentage of inputs that successfully lead LLMs to generate harmful outputs. An output is determined to be harmful if it does not match with rejection phrases, such as “I’m sorry”, “I apologize” and “As an”. This is not a perfect measurement but works relatively well in practice since LLMs are trained to reject harmful replies. It is also easy to measure and interpret. For prompt learning methods tested on other tasks, we employ Accuracy (Acc) as the metric. The processing time is determined as the average time used for each iteration across all input samples and all iterations. In all experiments, we use 1 NVIDIA A100 GPU with 80GB memory unless mentioned otherwise.
@@
To determine the hyperparameters for probe sampling, including probe set size $k$, filtered set size reduction hyperparameter $R$, we construct a validation set of size $100$ from AdvBench by random sampling in the $400$ instances different from the test set. We follow  Zou et al. (2023) for the hyperparameters used in the original algorithm such as the size of the candidate set $B$. We provide detailed analysis of hyperparameters in Section 3.4. When we combine probe sampling with simulated annealing, we follow the same procedure to select hyperparameters. We use the same number of optimization steps $500$ as in GCG throughout the paper.
@@ 3.4 Further analysis
Here we also experiment with bigger draft models, some of which is of similar size to Llama2. We experiment with GPT-Neo (Gao et al., 2020), Flan-T5-base (Chung et al., 2022), BART (Lewis et al., 2019), Phi-1.5 (Li et al., 2023), TinyLlama (Zhang et al., 2024) and Sheared-LLaMA (Xia et al., 2023). Among them, Sheared-LLaMA might be the closest to Llama2 since it is a pruned version of Llama2. For TinyLlama, Phi and Sheared-LLaMA, we use 2 A100s with 80GB memory to fit the whole computation.
@@ 4 Related Work
In the field of acceleration, speculative sampling (Chen et al., 2023; Leviathan et al., 2023) is the most relevant to our method. They also use a draft model but its design cannot be directly applied to accelerate the GCG algorithm. REST (He et al., 2023) adopts the concept of speculative sampling but uses a retrieval approach based on a Trie to construct the candidate. The attention module has also been a focus of acceleration because of its quadratic nature (Dao et al., 2022; Cai et al., 2024). There have also been continuous interests in more efficient versions of Transformers (So et al., 2019; Dai et al., 2021; Liu et al., 2021; Gu et al., 2020, 2021). These architectural changes are complementary to our algorithm design and we leave it to future work to study their effects on the GCG algorithm.
LLM Jailbreaks.
LLM Jailbreaks have received considerable interests recently since due to the implications of applying LLMs widely in human society. Although there is a continuous effort to build safe and reliable LLMs, bypassing the safety mechanism of LLMs is not uncommon. For example, fine-tuning a safe LLM on a few data instances can easily breaks its safety guarantees (Qi et al., 2023; Lermen et al., 2023). Treating the jailbreak as a prompt optimization problem has also led to a certain level of success (Zou et al., 2023; Mökander et al., 2023; Liu et al., 2023a; Chao et al., 2023; Geisler et al., 2024). In addition, conversing in a ciphered language (Yuan et al., 2023), planting a backdoor during RLHF (Rando and Tramèr, 2023), using a less well-aligned language (Deng et al., 2023) and multi-modality (Shayegani et al., 2023) can also lead to successful jailbreaks. Researchers have also constructed large dataset of manual jailbreak prompts (Toyer et al., 2023).
Alignment of LLMs.
To build safe LLMs, alignments has also been a widely studied topic in the community (Stiennon et al., 2020; Ouyang et al., 2022). Efforts have been put into improving helpfulness (Cheng et al., 2023; Bai et al., 2022a), honesty (Kaddour et al., 2023; Liu et al., 2023b; Park et al., 2023; Xu et al., 2023), and harmlessness (Hartvigsen et al., 2022). Among these works, there has been a growing interest in using feedback from a LLM to perform alignment (Bai et al., 2022b; Gulcehre et al., 2023; Yuan et al., 2024; Burns et al., 2023). In particular, (Burns et al., 2023) studies the case where a stronger LLM uses feedback from a weaker LLM. Despite all the efforts, there has not been a definitive answer for LLM safety alignments, which also motivates our research in LLM safety.
@@ 5 Conclusion
In this paper, we propose an algorithm probe sampling that can effectively accelerate the GCG algorithm. We achieve an acceleration ranging from 2.1×2.1$\times$2.1 × to 6.3×6.3$\times$6.3 × in different scenarios on AdvBench. We illustrate the intuition and how the algorithm works through extensive experiments. Furthermore, this approach is also applied to general prompt optimization methods and other jailbreak techniques, including AutoPrompt, APE, and AutoDAN.
We believe the idea of using the probe agreement score to perform adaptive computation can be applied to cases other than GCG. For example, it could potentially be used to perform conditional computation for attention. Another direction is to extend the framework to the multi-modality case which can be interesting given the vast amount of video data. It would also be interesting to run a small draft model on the scale of web data to detect the existence of natural adversarial prompts.

### v3

@@ Abstract
Safety of Large Language Models (LLMs) has become a critical issue given their rapid progresses. Greedy Coordinate Gradient (GCG) is shown to be effective in constructing adversarial prompts to break the aligned LLMs, but optimization of GCG is time-consuming. To reduce the time cost of GCG and enable more comprehensive studies of LLM safety, in this work, we study a new algorithm called Probe sampling. At the core of the algorithm is a mechanism that dynamically determines how similar a smaller draft model’s predictions are to the target model’s predictions for prompt candidates. When the target model is similar to the draft model, we rely heavily on the draft model to filter out a large number of potential prompt candidates. Probe sampling achieves up to $5.6$ times speedup using Llama2-7b-chat and leads to equal or improved attack success rate (ASR) on the AdvBench. Furthermore, probe sampling is also able to accelerate other prompt optimization techniques and adversarial methods, leading to acceleration of 1.8×1.8$\times$1.8 × for AutoPrompt, 2.4×2.4$\times$2.4 × for APE and 2.4×2.4$\times$2.4 × for AutoDAN. [^1] Our code is publicly available at https://github.com/zhaoyiran924/Probe-Sampling.
@@ 1 Introduction
Ensuring the safety of Large Language Models (LLMs)  (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023; Jiang et al., 2023) has become a central theme of research. Despite continuous efforts, LLMs are prone to generate objectionable contents in various scenarios including using an adversarial suffix (Zou et al., 2023), further finetuning (Qi et al., 2024; Lermen and Rogers-Smith, 2024), ciphering (Yuan et al., 2024b) and multilingual settings (Deng et al., 2024). Among effective LLM adversarial attack works, Greedy Coordinate Gradient (GCG) (Zou et al., 2023) present a general and universal method as briefly illustrated in Figure 1.
@@
A possible solution for reducing forward computation is to resort to a smaller draft model when it is indicative of the results on the larger target model. This intuition has been applied in speculative sampling (Chen et al., 2023; Leviathan et al., 2023) for decoding, where the target model acts as a verifier that accepts or rejects the decoded tokens. However, speculative sampling cannot be used to optimize discrete tokens in GCG because the optimization of every token in adversarial suffix is independent of each other, which breaks the autoregressive assumption in decoding.
Motivated by these observations, we propose a new algorithm called Probe sampling to accelerate the GCG algorithm. Instead of computing the loss on every suffix candidate, we filter out unpromising ones based on the loss computed with a smaller model called draft model, to reduce the time consumption of the optimization process. Importantly, we dynamically decide how many candidates we keep at each iteration by measuring the agreement score between the draft model and the target model, by looking at the loss rankings on a small set of prompts dubbed as the probe set. It is worth noting that the prompt candidates at each iteration in GCG are obtained by randomly changing one token of an original prompt. As a result, the agreement score is adaptive to the original prompt. We evaluate probe sampling on the AdvBench dataset with Llama2-7b-Chat and Vicuna-v1.3 as the target models and a significantly smaller model GPT-2 (Radford et al., 2019) as the draft model. Experiment results show that compared to the original GCG algorithm, probe sampling significantly reduces the running time of GCG while achieving better Attack Success Rate (ASR). Specifically, with Llama2-7b-Chat, probe sampling achieves $3.5$ times speedup and an improved ASR of $81.0$ compared to GCG with $69.0$ ASR. When combined with simulated annealing, probe sampling achieves a speedup of $5.6$ times with a better ASR of $74.0$.
Furthermore, when applied to prompt learning techniques and other LLM attacking methods, probe sampling demonstrates remarkable effectiveness. Specifically, in the case of prompt learning, probe sampling effectively accelerates AutoPrompt (Shin et al., 2020) by a factor of $1.8$. Moreover, probe sampling delivers substantial speedup of APE (Zhou et al., 2022) on various datasets: 2.3×2.3$\times$2.3 × on GSM8K, 1.8×1.8$\times$1.8 × on MMLU and 3.0×3.0$\times$3.0 × on BBH. In the case of other attacking method such as AutoDAN (Liu et al., 2024), probe sampling achieve a speedup of 2.3×2.3$\times$2.3 × and 2.5×2.5$\times$2.5 × on AutoDAN-GA and AutoDAN-HGA respectively.
@@ 2.1 Background: Greedy Coordinate Gradient
where $x$ is a prompt that contains a harmful user query such as “Tell me how to build a bomb”, $y$ is the target sentence “Sure, here is how to build a bomb”, and $s$ is the adversarial suffix that is optimized to induce the generation of $y$. $p$ is the probability of a sentence output by a LLM. This objective can be decomposed into the summation of the negative log likelihood of individual tokens in the target sentence like a typical language modeling objective. $s$ is set to be a fixed length string in the GCG algorithm.
The optimization of the adversarial suffix $s$ is a non-trivial problem. Prior works (Guo et al., 2021; Wen et al., 2024) based on Gumbel-Softmax (Jang et al., 2016; Maddison et al., 2022) and soft prompt tuning (Lester et al., 2021) have achieved limited success, probably because the LLMs are well-aligned and the exceptionally large models magnifies the difference between a discrete choice and its continuous relaxations.
@@ 2.2 Probe Sampling
After obtaining the agreement score, we keep $(1-\alpha)*B/R$ candidates where $(1-\alpha)*B$ means that the filtered set size is a scale-down of the previous batch size $B$ and $R$ is a hyperparameter that determines a further scale down. When $\alpha$ is close to 00, meaning little agreement between the two models, we will use a filtered set size of $B/R$. When $\alpha$ goes to $1$, we almost filter out most of the candidates. With the filtered size determined, we can readily rank the candidates according to the draft model and filter the ones with higher losses. Finally, we evaluate the final loss on the filtered set using the target model and select the best candidate.
@@ 2.3 Applying Probe Sampling to Other Prompt Optimization Methods
Although prompt sampling was designed to accelerate GCG, the general idea of reducing forward computation can be applied on other prompt optimization methods, where there is usually a process of sampling prompt candidates and evaluating their performances. To see whether probe sampling can effectively accelerate other methods, we also apply probe sampling to two prompt learning methods AutoPrompt (Shin et al., 2020) and APE (Zhou et al., 2022). In addition, we apply probe sampling on AutoDAN (Liu et al., 2024), a genetic algorithm that can find natural jailbreak prompts.
@@ 3.1 Experiment Details
Following (Zou et al., 2023), we use Attack Success Rate (ASR) as the evaluation metric for GCG and AutoDAN, which is defined as the percentage of inputs that successfully lead LLMs to generate harmful outputs. An output is determined to be harmful if it does not match with rejection phrases, such as “I’m sorry”, “I apologize” and “As an”. This is not a perfect measurement but works relatively well in practice since LLMs are trained to reject harmful replies. It is also easy to measure and interpret. For prompt learning methods tested on other tasks, we employ Accuracy (Acc) as the metric. The processing time is determined as the average time used for each iteration across all input samples and all iterations. In all experiments, we use 1 NVIDIA A100 GPU with 80GB memory unless mentioned otherwise.
@@
To determine the hyperparameters for probe sampling, including probe set size $k$, filtered set size reduction hyperparameter $R$, we construct a validation set of size $100$ from AdvBench by random sampling in the $400$ instances different from the test set. We follow  (Zou et al., 2023) for the hyperparameters used in the original algorithm such as the size of the candidate set $B$. We provide detailed analysis of hyperparameters in Section 3.4. When we combine probe sampling with simulated annealing, we follow the same procedure to select hyperparameters. We use the same number of optimization steps $500$ as in GCG throughout the paper.
@@ 3.2 Main Results
Table 2: Transferability of Probe sampling with different draft models.
Method Direct Transfer
Llama2-7b Vicuna-7b Mistral-7b
GCG $69.0$ $89.0$ $\mathbf{86.0}$
PS (GPT-2) $85.0$ $92.0$ $83.0$
PS (ShearedLlaMa) $\mathbf{91.0}$ $\mathbf{93.0}$ $85.0$
PS (Flan-T5) $57.0$ $78.0$ $69.0$
Table 3: Transferability of Probe sampling with different filtered set size $(1-\alpha)*B/R$.
Method Direct Transfer
Llama2-7b Vicuna-7b Mistral-7b
GCG $69.0$ $89.0$ $\mathbf{86.0}$
PS ($R=64$) $60.0$ $77.0$ $74.0$
PS ($R=8$) $\mathbf{85.0}$ $\mathbf{92.0}$ $83.0$
PS ($R=1$) $79.0$ $88.0$ $84.0$
@@
Transferability
Table 3 shows probe sampling’s transferability across draft models based on Llama2-7b-Chat to various target models. We find that it maintains transferability when using draft models like GPT-2 and SheardLlaMa, which preserve the original ASR of plain GCG. However, draft models that significantly degrade initial performance, such as Flan-T5, impair transferability. Table 3 examines probe sampling transferability across filtered set sizes. Results align with prior findings: probe sampling minimally impacts transferability with appropriate parameters but decreases performance when Llama2-7b-chat’s direct ASR is low such as $R=64$.
@@ 3.4 Further analysis
Here we also experiment with bigger draft models, some of which is of similar size to Llama2. We experiment with GPT-Neo (Gao et al., 2020), Flan-T5-base (Chung et al., 2024), BART (Lewis et al., 2019), Phi-1.5 (Li et al., 2023), TinyLlama (Zhang et al., 2024) and Sheared-LLaMA (Xia et al., 2023). Among them, Sheared-LLaMA might be the closest to Llama2 since it is a pruned version of Llama2. For TinyLlama, Phi and Sheared-LLaMA, we use 2 A100s with 80GB memory to fit the whole computation.
@@ 4 Related Work
Alignment of LLMs.
To build safe LLMs, alignments has also been a widely studied topic in the community (Stiennon et al., 2020; Ouyang et al., 2022). Efforts have been put into improving helpfulness (Bai et al., 2022a; Cheng et al., 2023), honesty (Kaddour et al., 2023; Liu et al., 2023; Xu et al., 2023), and harmlessness (Hartvigsen et al., 2022). Among these works, there has been a growing interest in using feedback from a LLM to perform alignment (Bai et al., 2022b; Gulcehre et al., 2023; Burns et al., 2024; Yuan et al., 2024a). Despite all the efforts, there has not been a definitive answer for LLM safety alignments, which also motivates our research in LLM safety.
Discrete Prompt Optimization.
Attacking LLMs via adversarial prompt can be formulated as a discrete prompt optimization problem (Zou et al., 2023). In this context, attacking algorithms strive to discover superior prompts that effectively steer aligned LLMs toward generating adversarial answers. Some approaches leverage LLMs themselves to iteratively refine prompts (Xu et al., 2022; Pryzant et al., 2023). However, aligned LLMs may resist refining adversarial prompts, rendering these methods ineffective. Other strategies employ RL-based prompt optimization techniques such as those in (Mingkai and Jianyu, 2022; Lu et al., 2023), necessitating additional MLP training with extensive adversarial data and specific reward design. Moreover, other models introduced in (Cho et al., 2023; Long et al., 2024) to help with prompt optimization must remain unaligned, particularly in jailbreak scenarios (Chao et al., ). However, their performance tends to be limited, especially when dealing with strongly fine-tuned models like Llama2-Chat.
LLM Jailbreaks.
LLM Jailbreaks have received considerable interests recently since due to the implications of applying LLMs widely in human society. Although there is a continuous effort to build safe and reliable LLMs, bypassing the safety mechanism of LLMs is not uncommon. For example, fine-tuning a safe LLM on a few data instances can easily breaks its safety guarantees (Qi et al., 2024; Lermen and Rogers-Smith, 2024). Treating the jailbreak as a prompt optimization problem has also led to a certain level of success (Zou et al., 2023; Mökander et al., 2023; Liu et al., 2024; Chao et al., ; Geisler et al., 2024). In addition, conversing in a ciphered language (Yuan et al., 2024b), planting a backdoor during RLHF (Rando and Tramèr, 2023), using a less well-aligned language (Deng et al., 2024) and multi-modality (Shayegani et al., 2024) can also lead to successful jailbreaks. Researchers also construct large dataset of manual jailbreak prompts (Toyer et al., 2023).
@@
In the field of acceleration, speculative sampling (Chen et al., 2023; Leviathan et al., 2023) is the most relevant to our method. They also use a draft model but its design cannot be directly applied to accelerate the GCG algorithm. REST (He et al., 2024) adopts the concept of speculative sampling but uses a retrieval approach based on a Trie to construct the candidate. The attention module has also been a focus of acceleration because of its quadratic nature (Dao et al., 2022; Cai et al., 2024). There have also been continuous interests in more efficient versions of Transformers (So et al., 2019; Dai et al., 2021; Liu et al., 2021; Gu et al., 2020, 2021). These architectural changes are complementary to our algorithm design and we leave it to future work.
@@ 5 Conclusion
In this paper, we propose an algorithm probe sampling that can effectively accelerate the GCG algorithm. We achieve an acceleration ranging from 2.1×2.1$\times$2.1 × to 6.3×6.3$\times$6.3 × in different scenarios on AdvBench. We illustrate the intuition and how the algorithm works through extensive experiments. Furthermore, this approach is also applied to general prompt optimization methods and other jailbreak techniques, including AutoPrompt, APE, and AutoDAN. We believe the idea of using the probe agreement score to perform adaptive computation can be applied to cases other than GCG. For example, it could potentially be used to perform conditional computation for attention. Another direction is to extend the framework to the multi-modality case which can be interesting given the vast amount of video data. It would also be interesting to run a small draft model on the scale of web data to detect the existence of natural adversarial prompts.
@@
Acknowledgements
This research is partially supported by the National Research Foundation Singapore under the AI Singapore Programme (AISG Award No: AISG2-TC-2023-010-SGIL) and the Singapore Ministry of Education Academic Research Fund Tier 1 (Award No: T1 251RES2207). Xuan Long Do is supported by the A*STAR Computing and Information Science (ACIS) scholarship. We thank Liwei Kang for insightful discussion, Liying Cheng for helping with plotting figures.